OpenStudy (anonymous):

math question help? The function f(t) = 5 cos(pi over 4t) + 11 represents the tide in Dark Sea. It has a maximum of 16 feet when time (t) is 0 and a minimum of 6 feet. The sea repeats this cycle every 8 hours. After six hours, how high is the tide? 11 feet 6 feet 8.5 feet 10.5 feet I think its 6 feet, am I right?

OpenStudy (alekos):

If you substitute t=6 hrs what do you get?

OpenStudy (anonymous):

-2?

OpenStudy (anonymous):

or =5.215?

OpenStudy (alekos):

no way. what's cos(π/4)t when t = 6?

OpenStudy (anonymous):

sqrt2/2 ?

OpenStudy (alekos):

no, no. it becomes cos(6π/4) which is?

OpenStudy (anonymous):

o?

OpenStudy (alekos):

that's right so what's the final answer in feet?

OpenStudy (anonymous):

so it is 6? @alekos

OpenStudy (alekos):

what's 5cos(3π/2) + 11 ? remember that cos(6π/4) = cos(3π/2) = 0